# Reduced-Order Modeling of Unsteady Fluid Flow Using Neural Network Ensembles

### v1

@@ 3.1 Lid-Driven Cavity
Figure  10 shows snapshots of $u$ and $v$ at $T=750000$ as well as ROM errors averaged over the last 10% of the simulation for a single seed. The errors are significantly lower throughout the computational domain when using LSTM ensembles, which are greatest in the wake of the cylinder and immediately downstream of it. Table 6 lists the seed-averaged relative errors in $u$ and $v$ at the test points. The ensemble ROM again offers better performance in predicting both fields at all of the test points. Table 7 lists the standard deviation of these errors, which are again lower at all of the test points. At the fourth test case index ($\bm{\mu^{*}}=[57,171.0]$), the errors and standard deviations are similar. This test point lies in the middle of the parameter space and is well-surrounded by training samples; as a result, the advantage gained using an ensemble method may be marginal. Table 8 lists the computational costs associated with the problem. Using the ROM for inference offers a speed-up of approximately 27x over CFD (again, the given CFD cost is the portion of the simulation over the prediction time horizon).
@@
Table 6: Average relative errors in $u$ and $v$ for the 2D cylinder case.
@@
Table 7: Standard deviation in relative errors of $u$ and $v$ for the 2D cylinder case.
@@
Table 8: Computational costs associated with the 2D cylinder case.
@@ 4 Conclusion
This work presents a fully data-driven ROM framework for time-dependent physics simulations using convolutional autoencoders and LSTM ensembles. The combined framework, referred to as the CAE-eLSTM ROM, is used to mitigate the issue of error propagation when performing autoregressive time-series prediction at unseen data sets, a common problem in data-driven CFD applications. The ROM obtains a low-dimensional solution manifold using convolutional autoencoders, a type of artificial neural network useful for reconstructing spatially distributed data. Using the encoder section of the autoencoder, the low-dimensional latent variables of the training data are used to generate multivariate time-series sequences that are used to train LSTMs, a type of recurrent neural network useful for modeling sequential data. Bagging, a popular ensemble learning technique, is used to train multiple base LSTMs on subsets of the training sequences sampled randomly with replacement. Ensemble learning is chosen as it is a useful tool for improving the stability and accuracy of machine learning methods. At unseen test points, the full-order model is run to obtain an initial sequence of latent variables through the encoder and autoregressive time-series prediction is used to predict the latent variables over a prescribed time horizon. The predicted latent variables are then passed through the decoder to obtain spatial reconstructions of the full-order model at different points in time.
@@
When applied to two incompressible, unsteady, laminar fluid dynamics problems, the proposed ROM exhibits high levels of accuracy when compared to a ROM using a single LSTM model. The use of LSTM ensembles significantly diminishes the error propagation issue, with latent variable trajectories from different weight initializations showing low levels of variance. Error metrics show that using the ensemble ROM leads to better predictive performance at all test points. Furthermore, the fully data-driven nature of the ROM allows it to be applied to two different CFD solvers. Although the cost associated with training ensemble neural networks is high, bagging allows for models to be trained in parallel, and multi-GPU architectures can be used to significantly reduce the cost. Future work will focus on further developing the method to apply it to three-dimensional geometries, problems exhibiting turbulent flow, and more complex geometries.
@@ Appendix B Convolutional autoencoder architectures
The convolutional autoencoder architectures for the lid-driven cavity and 2D cylinder cases are given in tables 9 and 10 respectively. Both architectures are similar and use convolutional layers in the encoder section to progressively reduce the number of pixels or grow the number of filters in subsequent layers. Convolutional layers are usually followed max-pooling layers that have batch normalization [38] applied to them (referred to as pool-norm layers). Batch normalization normalizes the input of each layer over a mini-batch, reducing internal covariate shift, leading to more efficient gradient flow during backpropagation. Fully connected layers are present in both the encoder and decoder. Although their inclusion leads to a larger number of network parameters, reconstruction accuracies are significantly improved. The decoder consists of convolutional transpose layers that are used to progressively increase the number of pixels. The output layer uses a sigmoid activation function to constrain the range of outputs to [0,1]. The leaky ReLU activation function is used for convolutional and fully connected layers, and is given as,
@@
Table 9: Convolutional autoencoder architecture for the lid-driven cavity case.
@@
Table 10: Convolutional autoencoder architecture for the 2D cylinder case.

### v2

@@ 3.1 Lid-Driven Cavity
Figure  11 shows snapshots of $u$ and $v$ at $T=750000$ as well as ROM errors averaged over the last 10% of the simulation for a single seed. The errors are significantly lower throughout the computational domain when using LSTM ensembles, which are greatest in the wake of the cylinder and immediately downstream of it. Table VI lists the seed-averaged relative errors in $u$ and $v$ at the test points. The ensemble ROM again offers better performance in predicting both fields at all of the test points. Table VII lists the standard deviation of these errors, which are again lower at all of the test points. At the fourth test case index ($\bm{\mu^{*}}=[57,171.0]$), the errors and standard deviations are similar. This test point lies in the middle of the parameter space and is well-surrounded by training samples; as a result, the advantage gained using an ensemble method may be marginal. Table VIII lists the computational costs associated with the problem. Using the ROM for inference offers a speed-up of approximately 27x over CFD (again, the given CFD and ROM inference costs are for the portion of the simulation over the prediction time horizon). Due to the increased domain size and amount of training data, the CAE and LSTM are more expensive to train when compared to the lid-driven cavity case.
@@
Table VI: Average relative errors in $u$ and $v$ for the 2D cylinder case. Boldface denotes the lowest values.
@@
Table VII: Standard deviation in relative errors of $u$ and $v$ for the 2D cylinder case. Boldface denotes the lowest values.
@@
Table VIII: Computational costs associated with the 2D cylinder case.
@@ 4 Conclusion
This work presents a fully data-driven ROM framework for time-dependent physics simulations using convolutional autoencoders and LSTM ensembles. The combined framework, referred to as the CAE-eLSTM ROM, is used to mitigate the issue of error propagation when performing autoregressive time-series prediction at unseen data sets, a common problem in data-driven CFD applications. The ROM obtains a low-dimensional solution manifold using convolutional autoencoders, a type of artificial neural network useful for reconstructing spatially distributed data. Using the encoder section of the autoencoder, the low-dimensional latent variables of the training data are used to generate multivariate time-series sequences that are used to train LSTMs, a type of recurrent neural network useful for modeling sequential data. Bagging, a popular ensemble learning technique, is used to train multiple base LSTMs on subsets of the training sequences sampled randomly with replacement. Ensemble learning is chosen as it is a useful tool for improving the stability and accuracy of machine learning methods. At unseen test points, the full-order model is run to obtain an initial sequence of latent variables through the encoder and autoregressive time-series prediction is used to predict the latent variables over a prescribed time horizon. The predicted latent variables are then passed through the decoder to obtain spatial reconstructions of the FOM at different points in time.
@@
When applied to two incompressible, unsteady, laminar fluid dynamics problems, the proposed ROM exhibits high levels of accuracy when compared to a ROM using a single LSTM model. The use of LSTM ensembles significantly diminishes the error propagation issue, with latent variable trajectories from different weight initializations showing low levels of variance. Error metrics show that using the ensemble ROM leads to better predictive performance at all test points. Furthermore, the fully data-driven nature of the ROM allows it to be applied to two different CFD solvers. Although the cost associated with training ensemble neural networks is high, bagging allows for models to be trained in parallel, and multi-GPU architectures can be used to significantly reduce the cost. The addition of an ensemble model for time-series predictions, while simple, greatly improves the performance of the ROM and does not require additional training data. A limitation of the presented framework is its significantly increased offline training cost. While this can be alleviated to a great degree by training weak learners in parallel, multi-GPU platforms are expensive and may not be readily available for practitioners. Additionally, the use of convolutional autoencoders limits the application of the ROM to structured grids unless a mesh interpolation method is used. Future work will focus on further developing the method to apply it to three-dimensional geometries, problems exhibiting turbulent flow, and more complex geometries involving unstructured meshes. Problems consisting of time-series datasets that exhibit non-smooth variations over long periods of time will be of particular interest, as they are more representative of real-world applications and can reap greater benefits from ensemble learning. Computationally efficient optimal hyperparameter selection methods will also be explored to lower the offline costs of training the ROM.
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Rakesh Halder: Conceptualization (equal), Formal Analysis (lead), Methodology (lead), Software (equal), Writing/Original Draft Preparation (lead)
Mohammadmehdi Ataei: Conceptualization (equal), Software (equal), Supervision (equal), Writing/Review & Editing (equal)
Hesam Salehipour: Conceptualization (equal), Software (equal), Supervision (equal), Writing/Review & Editing (equal)
Krzysztof Fidkowski: Writing/Review & Editing (equal)
Kevin Maki: Writing/Review & Editing (equal)
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@ Appendix B Convolutional autoencoder architectures
The convolutional autoencoder architectures for the lid-driven cavity and 2D cylinder cases are given in Tables IX and X respectively. Both architectures are similar and use convolutional layers in the encoder section to progressively reduce the number of pixels or grow the number of filters in subsequent layers. Convolutional layers are usually followed max-pooling layers that have batch normalization [46] applied to them (referred to as pool-norm layers). Batch normalization normalizes the input of each layer over a mini-batch, reducing internal covariate shift, leading to more efficient gradient flow during backpropagation. Fully connected layers are present in both the encoder and decoder. Although their inclusion leads to a larger number of network parameters, reconstruction accuracies are significantly improved. The decoder consists of convolutional transpose layers that are used to progressively increase the number of pixels. The output layer uses a sigmoid activation function to constrain the range of outputs to [0,1]. The leaky ReLU activation function is used for convolutional and fully connected layers, and is given as,
@@
Table IX: Convolutional autoencoder architecture for the lid-driven cavity case.
@@
Table X: Convolutional autoencoder architecture for the 2D cylinder case.